2017 December

From EOVSA Wiki
Jump to: navigation, search

Back to Observing Log

Dec 13

02:16 UT This morning (Dec. 12) the DPP was up, but showed signs of having been rebooted somehow, so recording was not working. At the same time, I found that the ROACH boards were not producing enough packets. I restarted everything, but at that point the delays were not correct. Tonight I analyzed the last calibration scan of the day, and now the delays should be okay for tomorrow. Another thing I noticed (not sure when this happened, but probably on Dec. 11) is that Ant 11 had the wrong crio.ini file, so I had to load the correct one.

15:21 UT I found ant5 had the wrong crio.ini. I loaded the correct one. I also see that the ACC is causing the glitches in the data again, so it has to be rebooted. Since I was too late before the refcal started, I will wait until 16:05 UT when we go to the Sun.